OpenStudy (anonymous):

a 2600 kg truck travelling at 72 km/h slams on the brakes and skids to a stop. The frictional force from the road is 8200 N. Use the relationship between kinetic energy and mechanical work to determine the distance it takes for the truck to stop.

sam (.sam.):

-fd=(1/2)m(v^2-u^2) \[-d=\frac{(1/2)m(v^2-u^2)}{F}\] \[d=-\frac{(1/2)(2600)(0^2-(20)^2)}{8200}=63.4meters\]